the sum of 2 consecutive even integers is 70. find the larger number. Please help me solve this, I need to the the process
Answer by nerdybill(7384)

the sum of 2 consecutive even integers is 70. find the larger number. Please help me solve this, I need to the the process
.
Let x = larger of two consecutive even integers
then
x-2 = smaller of two consecutive even integers
.
from: "the sum of 2 consecutive even integers is 70" we get our equation:
x + (x-2) = 70
x + x - 2 = 70
2x - 2 = 70
2x = 72
x = 36 (answer)